OpenStudy (anonymous):

find the y-intercept of the line through the point (-1, 10/3) with slope -1.

OpenStudy (lucaz):

write the equation of this line first

OpenStudy (ybarrap):

$$ y=mx+b $$ m is your slope b is your y-intercept We know the point <-1,10/3> and the slope, m=-1, so $$ \cfrac{10}{3}=(-1)(-1) + b\\ b=\cfrac{10}{3}-1=\cfrac{10-3}{3}=\cfrac{7}{3} $$ So $$ y=mx+b\\ =-x+\cfrac{7}{3} $$ This is in y-intercept form. Make sense?